# Cosmological singularity, conformal anomaly and symmetric polynomials

### v1

@@ Abstract
We consider a spacetime singularity at $t=0$ arising in a Kasner-type metric that solves the gravitational equations modified by quantum effects of a conformal field theory (CFT). The resulting constraints can be solved efficiently when expressed in terms of symmetric polynomials. Focusing first on the trace part of the modified gravitational equation, we determine the corresponding solution surfaces in Kasner-parameter space. The geometry of these surfaces depends sensitively on the ratio $\eta=A/C$, the quotient of the conformal charges characterizing the underlying CFT. We then fully integrate the conformal anomaly near the singularity for a generic Kasner-type metric and obtain the corresponding stress-energy tensor. Its components are expressed in terms of three symmetric polynomials (of degrees $2$, 33 and $4$) and depend on seven arbitrary constants, which may be interpreted as parameterizing different choices of the quantum state at the singularity. By imposing a set of constraints we reduce this parameter space to a single free constant. Subsequently, we solve, at leading order near the singularity, the modified gravitational equations. Among the admissible solutions, we identify, in particular, those that develop a curvature singularity while remaining geodesically complete.
@@ 1 Introduction
A major step toward such an understanding was achieved in the seminal works of Belinsky, Khalatnikov and Lifshitz (BKL) [2], who proposed a detailed asymptotic description of the metric near a spacetime singularity. According to the BKL scenario, temporal derivatives dominate over spatial ones as the singularity is approached, reducing the Einstein equations to a system of ordinary differential equations. In this regime, the metric evolves through a sequence of anisotropic contraction phases, the Kasner epochs, each characterized by a set of Kasner exponents and separated by abrupt transitions or bounces.The resulting oscillatory, and in fact chaotic, sequence of epochs is one of the most striking manifestations of nonlinear behaviour in general relativity. Further progress in the study of this chaotic dynamics was made in [3].
@@
In this paper, we systematically analyse the backreacted semiclassical geometry within the class of Kasner-type metrics. We identify the constraints on the Kasner parameters arising from the conformal anomaly. We then integrate the anomaly explicitly in this class of backgrounds, allowing us to extract the full set of tensorial constraints. These constraints can be solved in closed form in terms of symmetric polynomials of the Kasner parameters.
@@ 2.3 Regularity conditions for singular metrics
2. Integral curvature is finite. Although the square of the Riemann tensor behaves as $R_{\alpha\beta\mu\nu}^{2}\sim t^{-4}$ and is therefore divergent at $t=0$, its integral $\int dt\,\sqrt{-g}R_{\alpha\beta\mu\nu}^{2}$ can nevertheless be finite, provided that
@@
Consider a light ray propagating along the $x^{1}$ direction. The change in the affine parameter between a point at finite $t$ and the singularity at $t=0$ behaves as $\Delta\lambda\sim\int t^{a}dt$. This integral diverges provided $a\leq-1$. Extending this argument to null rays propagating along any spatial direction $x^{i}\,i=1,2,3$, we obtain the more general condition,
@@
Geodesically complete spacetime with curvature singularity. This analysis shows that for certain negative values of the parameters, the spacetime near the singularity attains infinite volume and its lightlike geodesics become complete. Since spacetime singularities are typically characterized by geodesic incompleteness, the case in which all parameters $(a,b,c)$ are negative and satisfy (2.9) is particularly appealing, as it appears to evade the usual pathological features.
@@ 2.4 Symmetric polynomials
Other symmetric polynomials of a fixed degree can be constructed from these three elementary ones. In degree 2 there are two independent elementary symmetric polynomials: $e_{1}^{2}$ and $e_{2}$. In degree 3 there are three polynomials: $e_{1}^{3}$, $e_{1}e_{2}$ and $e_{3}$. In degree 4 there are four: $e_{1}^{4}$, $e_{1}^{2}e_{2}$, $e_{2}^{2}$, $e_{1}e_{3}$. Below we provide some useful identities:
@@
It is often convenient, and in some cases particularly useful, to regard the symmetric polynomials $(e_{1},e_{2},e_{3})$ as independent variables in place of the parameters $(a,b,c)$. The Jacobian of this change of variables is proportional to (ab)(bc)(ac)(a-b)(b-c)(a-c), and is therefore non-vanishing whenever the three parameters $(a,b,c)$ are distinct.
Consistency bounds on symmetric polynomials. One should stress that, although a priori parameters $(a,b,c)$ may take any real (positive or negative) values, it is not the case for the symmetric polynomials $(e_{1},e_{2},e_{3})$. There appears a number of bounds based on various mathematical inequalities, see [14]. One restriction comes from the relation
@@
On the other hand, the AM-GM inequality implies that for any non-negative $a,b,c$ holds inequality
@@ 2.5 Classical Kasner metric
In the classical case [12] the parameters $(a,b,c)$ are those for which metric (2.1) solves the Einstein equations in vacuum, $G^{\mu}_{\nu}=0$. The equation $G^{0}_{0}=0$ implies that the respective symmetric polynomial $e_{2}=0$. The equation $G^{1}_{1}$ explicitly separates coordinate $x_{1}$ from $x_{2}$ and $x_{3}$ and hence the respective equation is symmetric under permutations of $b$ and $c$ but not under permutations that involve parameter $a$. Indeed, as is seen from (2.2), equation $G^{1}_{1}=0$ implies that (taking that $e_{2}=0$)
@@
Geometrically, in the space of parameters $(a,b,c)$, it is represented by a circle that is intersection of a sphere of unite radius ($a^{2}+b^{2}+c^{2}=1$) and a plane ($a+b+c=1$). The entire circle corresponds to all possible (positive and negative) values of parameter $u$.
Alternatively, one can use the symmetric polynomials $(e_{1},e_{2},e_{3})$ as independent variables. As we have seen already the symmetric polynomials $e_{1}=1$ and $e_{2}=0$. The polynomial $e_{3}$ takes negative values in the interval $\frac{4}{27}\leq e_{3}\leq 0$. The extremal negative value is reached for values $(a=-\frac{1}{3},b=c=\frac{2}{3})$. So that in the space of variables $(e_{1},e_{2},e_{3})$ the solution is given by an interval $(e_{1}=1,e_{2}=0,e_{3}=v\,,\,-\frac{4}{27}\leq v\leq 0)$.
@@ 3.1 General set-up
where $\kappa=8\pi G$. For free quantum fields $W_{Q}=-\frac{1}{2}\sum_{s}(-1)^{s}\det\mathop{\rm ln}\nolimits\Box_{(s)}$, where $s$ is spin and $\Box_{(s)}$ is the respective field operator. Generally, $W_{Q}$ is a rather complicated non-local functional of the background metric. In the case when the quantum fields are conformal an important local quantity that characterises $W_{Q}$ is available: the conformal anomaly ${\cal A}=g^{\mu\nu}\frac{\delta W_{Q}}{\delta g^{\mu\nu}}$. That is why in the present paper we are focused on the case of conformal quantum field theory.
Variation of the gravitational action with respect to metric gives us the modified Einstein equations,
@@
We stress that the background metric is not fixed but appears as a solution to this equation. We call it backreacted metric as it includes the backreaction of the quantum fields on the background metric. The quantum stress-energy tensor $T^{Q}_{\mu\nu}$ is covariantly conserved,
@@
As we have already mentioned, $T^{Q}_{\mu\nu}$ is generally unknown and it is defined up to a number of ambiguities. However, its trace is a known local function of metric. So that the trace of the gravitational equations (3.2) is local and unambiguous,
@@
In the present section we shall analyze this equation for the anisotropic metric whose near-singularity behaviour is given by (2.1).
@@ 3.2 Master equation
First, we compute $W^{2}$ and $E_{4}$ near the singularity. For a Kasner type metric (2.1) one finds that
@@ 3.3 Some particular cases
We now consider some particular in which the Master equation is simplified.
Isotropic case: $a=b=c$. When all three parameters coincide the Weyl squared vanishes, $W^{2}=0$, so that the anomaly is determined only the the Euler term:
@@
For values $\eta>1$ there appears a closed curve component in the domain $a<0,b>0$. This curve approaches quite closely the origin but does not in fact reaches it remaining always in the region $a<0,b>0$.
For negative values $\eta<0$ the curve has three components: one open component in domain $a<0,b>0$, the other open component in domain $a>0,b<0$ and a closed curve component in the domain $a>0,b>0$ that passes by the points $(a=0,b=0)$ and $(a=1,b=0)$. These two points correspond to spacetime without a singularity.
@@ 3.4 General shape of the Master equation surfaces
In a generic case solution to the Master equation defines a two-dimensional surface in the space of parameters $(a,b,c)$. The shape of this surface crucially depends on the parameter $\eta=A/C$ that is determined uniquely by the quantum conformal field theory in question. For unitary CFT parameter $\eta$ is positive and satisfies certain unitarity bounds that were studied in the literature. As we already explained this above in the present study we do not restrict ourselves to only unitary theories and thus prefer to explore all possible values of $\eta$ taking all possible (positive or negative) real numbers. In Figures 4-6 we present some typical situations with the representative values of $\eta$.
The analysis shows that from $\eta$ changing in the interval $0<\eta<1$ the surface contains of two open shells. One of them lies in the quadrant of positive values of parameters $(a,b,c)$ and bounded by plane $a+b+c=3$. So that for all points in this branch one has that $a+b+c>3$. The respective spacetime has regular integral Riemann curvature, see (2.7).
The second open shell lies in the region of negative values of all parameters $(a,b,c)$. It contains a conical point that has coordinates $(0,0,0)$. It has three open infinite domains where the conditions (2.9) to have the complete geodesics are satisfied. The respective spacetimes have infinite 4-volume near the singularity.
The entire surface is presented in Figure 4 for value $\eta=1/2$. Note that the points where all parameters vanish and where one of parameters is $1$ and the others vanish are not shown since they lie outside the continuous surface and represent the isolated points in the parameters space.
For values $\eta>1$ the surface still has two open shells as in the case above. However, there also appears three closed identical shells each of them lying in a region where one of the parameters is negative and the two others are positive. A surface of this type is shown in Figure 5, where in particular the closed shell is shown in the region $(a<0,b>0,c>0)$. The closed shell then passes by the points $(a=0,b=1,c=0)$ and $(a=0,b=0,c=1)$ that now are not isolated points and are continuously related to other points in the closed shell component. Geometrically these three closed shells are identical due to the permutation symmetry in the parameter space $(a,b,c)$.
@@
Summarizing, the configurations, that are similar to the classical Kasner solution, where one of the parameters is negative and the other two are positive, occur only for $\eta>1$ and $\eta<0$, and are absent for $0<\eta<1$. For $\eta>1$, the corresponding points lie on a compact surface, so the Kasner parameters remain bounded. In contrast, for $\eta<0$, the corresponding surface is non-compact, and the absolute values of the parameters are unbounded.
@@ 3.5 Solving the Master equation in terms of symmetric polynomials
The solution (3.12) has two branches depending on the sign. The expression under the square root is non-negative for two intervals: $e_{1}\geq 3$ and $e_{1}\leq\frac{1}{3}$. For $e_{1}=0$ in the positive sign branch $e^{-}_{2}=0$. This is the only point when $e_{2}$ vanishes. These values ($e_{1}=e_{2}=0$) in terms of parameters $(a,b,c)$ has the only solution $(0,0,0)$. It is the spacetime without a singularity.
@@
The bound (2.13) imposes extra constraint on possible values of the symmetric polynomials. One finds that in the interval $e_{1}\geq 3$ the bound selects only the negative branch, $e^{-}_{2}$. In the interval $0\leq e_{1}\leq\frac{1}{3}$ both branches violate the bound while in the interval $e_{1}\leq 0$ the positive sign branch satisfies the bound. This is shown in Figure 7.
@@
Note, that this solution saturates the bound (2.13). It is possible only if all parameters $a,b,c$ are equal, $a=b=c=1$. This solution we have already mentioned above. It corresponds to a symmetric spacetime with a singularity at $t=0$.
@@
4 Integrating conformal anomaly
In the preceding sections, we analyzed the trace part of the modified gravitational equations that incorporate the conformal anomaly. The full gravitational equations, however, contain additional components that introduce further constraints. Geometrically, these constraints select a subspace of solutions within the surface ${\cal M}(a,b,c)=0$ considered earlier. To investigate their implications, one must first determine the general structure of the anomalous stress-energy tensor. In this section, we examine the complete set of equations for the stress-energy tensor $T_{\mu\nu}=\bigl(T_{00}(t),T_{11}(t),T_{22}(t),T_{33}(t)\bigr)$ of the quantum CFT and provide a full integration of these equations.

### v2

@@ Abstract
We consider a spacetime singularity at $t=0$ arising in a Kasner-type metric that solves the gravitational equations modified by quantum effects of a conformal field theory (CFT). The resulting constraints can be solved efficiently when expressed in terms of symmetric polynomials. Focusing first on the trace part of the modified gravitational equation, we determine the corresponding solution surfaces in the Kasner parameter space. The geometry of these surfaces depends sensitively on the ratio $\eta=A/C$, the quotient of the conformal charges characterizing the underlying CFT. We then fully integrate the conformal anomaly near the singularity for a generic Kasner-type metric and obtain the corresponding stress-energy tensor. Its components are expressed in terms of three symmetric polynomials (of degrees $2$, 33 and $4$) and depend on seven arbitrary constants, which may be interpreted as parameterizing different choices of the quantum state at the singularity. By imposing a set of constraints we reduce this parameter space to a single free constant. Subsequently, we solve, at leading order near the singularity, the modified gravitational equations. Among the admissible solutions, we identify, in particular, those that develop a curvature singularity while remaining geodesically complete.
@@ 1 Introduction
A major step toward such an understanding was achieved in the seminal works of Belinsky, Khalatnikov and Lifshitz (BKL) [2], who proposed a detailed asymptotic description of the metric near a spacetime singularity. According to the BKL scenario, time derivatives dominate over spatial ones as the singularity is approached, reducing the Einstein equations to a system of ordinary differential equations. In this regime, the metric evolves through a sequence of anisotropic contraction phases, the Kasner epochs, each characterized by a set of Kasner exponents and separated by abrupt transitions or bounces.The resulting oscillatory, and in fact chaotic, sequence of epochs is one of the most striking manifestations of nonlinear behaviour in general relativity. Further progress in the study of this chaotic dynamics was made in [3].
@@
In this paper, we systematically analyze the backreacted semiclassical geometry within the class of Kasner-type metrics. We identify the constraints on the Kasner parameters arising from the conformal anomaly. We then integrate the anomaly explicitly in this class of backgrounds, which allows us to extract the full set of tensorial constraints. These constraints can be solved in closed form in terms of symmetric polynomials of the Kasner parameters.
@@ 2.3 Regularity conditions for singular metrics
2. Integrated curvature is finite. Although the square of the Riemann tensor behaves as $R_{\alpha\beta\mu\nu}^{2}\sim t^{-4}$ and is therefore divergent at $t=0$, its integral $\int dt\,\sqrt{-g}R_{\alpha\beta\mu\nu}^{2}$ can nevertheless be finite, provided that
@@
Consider a light ray propagating along the $x^{1}$ direction. The change in the affine parameter between a point at finite $t$ and the singularity at $t=0$ behaves as $\Delta\lambda\sim\int t^{a}dt$. This integral diverges when $a\leq-1$. Extending this argument to null rays propagating along any spatial direction $x^{i}\,i=1,2,3$, we obtain the more general condition,
@@
Geodesically complete spacetimes with curvature singularities. This analysis shows that for certain negative values of the parameters, the spacetime near the singularity attains infinite volume and its lightlike geodesics become complete. Since spacetime singularities are typically characterized by geodesic incompleteness, the case in which all parameters $(a,b,c)$ are negative and satisfy (2.9) is particularly appealing, as it appears to evade the usual pathological features.
@@ 2.4 Symmetric polynomials
Other symmetric polynomials of a fixed degree can be constructed from these. In degree 2 there are two independent elementary symmetric polynomials: $e_{1}^{2}$ and $e_{2}$. In degree 3 there are three polynomials: $e_{1}^{3}$, $e_{1}e_{2}$ and $e_{3}$. In degree 4 there are four: $e_{1}^{4}$, $e_{1}^{2}e_{2}$, $e_{2}^{2}$, $e_{1}e_{3}$. Below we provide some useful identities:
@@
It is often convenient, and in some cases particularly useful, to regard the symmetric polynomials $(e_{1},e_{2},e_{3})$ as independent variables instead of the parameters $(a,b,c)$. The Jacobian of this change of variables is proportional to (ab)(bc)(ac)(a-b)(b-c)(a-c), and is therefore non-vanishing whenever the three parameters $(a,b,c)$ are distinct.
Consistency bounds on symmetric polynomials. One should stress that, although a priori parameters $(a,b,c)$ may take any real (positive or negative) values, it is not the case for the symmetric polynomials $(e_{1},e_{2},e_{3})$. These satisfy a number of bounds based on various mathematical inequalities, see [14]. One restriction comes from the relation
@@
On the other hand, the inequality of arithmetic and geometric means (known as the AM-GM inequality) implies that for any non-negative $a,b,c$ the following inequality holds
@@ 2.5 Classical Kasner metric
In the classical case [12] the parameters $(a,b,c)$ are those for which the metric (2.1) solves the Einstein equations in vacuum, $G^{\mu}_{\nu}=0$. The equation $G^{0}_{0}=0$ implies that the corresponding symmetric polynomial $e_{2}=0$. The equation $G^{1}_{1}$ explicitly separates coordinate $x_{1}$ from $x_{2}$ and $x_{3}$ and hence the corresponding equation is symmetric under permutations of $b$ and $c$ but not under permutations that involve parameter $a$. Indeed, as is seen from (2.2), equation $G^{1}_{1}=0$ implies that (taking into account that $e_{2}=0$)
@@
Geometrically, in the space of parameters $(a,b,c)$, it is represented by a circle that is the intersection of a sphere of unit radius ($a^{2}+b^{2}+c^{2}=1$) and a plane ($a+b+c=1$). The entire circle corresponds to all possible values of the parameter $u$.
Alternatively, one can use the symmetric polynomials $(e_{1},e_{2},e_{3})$ as independent variables. As we have seen already the symmetric polynomials $e_{1}=1$ and $e_{2}=0$. The polynomial $e_{3}$ takes negative values in the interval $\frac{4}{27}\leq e_{3}\leq 0$. The lower bound is reached for values $(a=-\frac{1}{3},b=c=\frac{2}{3})$. So that in the space of variables $(e_{1},e_{2},e_{3})$ the solution is given by an interval $(e_{1}=1,e_{2}=0,e_{3}=v\,,\,-\frac{4}{27}\leq v\leq 0)$.
@@ 3.1 General set-up
where $\kappa=8\pi G$. For free quantum fields $W_{Q}=-\frac{1}{2}\sum_{s}(-1)^{s}\det\mathop{\rm ln}\nolimits\Box_{(s)}$, where $s$ is the spin and $\Box_{(s)}$ is the respective field operator. Generally, $W_{Q}$ is a rather complicated non-local functional of the background metric. In the case when the quantum fields are conformal an important local quantity that characterises $W_{Q}$ is available: the conformal anomaly ${\cal A}=g^{\mu\nu}\frac{\delta W_{Q}}{\delta g^{\mu\nu}}$. That is why in the present paper we focus on the case of conformal quantum field theory.
Variation of the gravitational action with respect to the metric gives us the modified Einstein equations,
@@
We stress that the background metric is not fixed but appears as a solution to this equation. We call it the backreacted metric as it includes the backreaction of the quantum fields on the background metric. The quantum stress-energy tensor $T^{Q}_{\mu\nu}$ is covariantly conserved,
@@
3.2 Our strategy
Our primary goal is to solve the semiclassical gravitational equations (3.2) for the class of Kasner-type metrics (2.1) and to determine the constraints on the parameters $(a,b,c)$. We achieve this goal in two steps.
Step 1. As already mentioned, the quantum stress tensor $T_{\mu\nu}$ is generally unknown and is defined up to a number of ambiguities. However, its trace is a known local function of the metric. Consequently, the trace of the gravitational equations (3.2) is local and unambiguous,
@@ 3.2 Our strategy
In the present section we shall analyze this equation for an anisotropic metric whose near-singularity behaviour is given by (2.1). To leading order, the left-hand side of equation (3.6) behaves as $1/t^{2}$ in the vicinity of the singularity, whereas the right-hand side behaves as $1/t^{4}$. Thus, to leading order in $1/t$ equation (3.6) for the metric in the general form (2.1) reduces to the requirement that the most divergent term on the right-hand side of (3.6) vanishes. This conditions leads to what we call the Master equation, which will be considered in the next subsection.
Proceeding with Step 1 has the advantage that it does not involve any ambiguities related to the integration of the continuity equation (3.3) for stress tensor $T^{Q}_{\mu\nu}$. However, the set of conditions on parameters $(a,b,c)$ obtained in Step 1 defines a rather broad surface of admissible values in the parameter space. In order to identify a more restricted subset of these conditions, we must consider the remaining tensorial components in equations (3.2). This is carried out in Step 2.
Step 2. We first solve explicitly the continuity equation (3.3) for the general form of the Kasner type metric (2.1), without imposing a priori any constraints on parameters $(a,b,c)$. This procedure introduces several constants that remain undetermined. Fixing these constants amounts to selecting a particular quantum state of the conformal field. This analysis is carried out in Sections 4 and 5, where we focus on the leading divergent terms behaving as $1/t^{4}$. The resulting stress tensor $T_{\mu\nu}$ is then substituted into equations (3.2). As in Step 1, the left-hand side of (3.2) behaves as $1/t^{2}$ while the right-hand side behaves as $1/t^{4}$. Therefore, solving equation (3.2) requires that the terms in $T^{\mu}_{\nu}$ diverging as $1/t^{4}$ vanish. This condition effectively reduces to the vanishing of three symmetric polynomials in the parameters $a,b,c$, as we show in Section 5. The solution of the resulting constraints is considered in Section 6.
@@ 3.3 Master equation
With these explanations, we proceed to solve the trace equation (3.6). First, we compute $W^{2}$ and $E_{4}$ near the singularity. For a Kasner type metric (2.1) one finds that
@@ 3.4 Some particular cases
We now consider some particular cases in which the Master equation simplifies.
Isotropic case: $a=b=c$. When all three parameters coincide the Weyl squared vanishes, $W^{2}=0$, so that the anomaly is determined only by the Euler term:
@@
For values $\eta>1$ there a closed curve component appears in the domain $a<0,b>0$. This curve approaches quite closely the origin but does not in fact reaches it remaining always in the region $a<0,b>0$.
For negative values $\eta<0$ the curve has three components: one open component in the domain $a<0,b>0$, another open component in the domain $a>0,b<0$ and a closed curve component in the domain $a>0,b>0$ that passes by the points $(a=0,b=0)$ and $(a=1,b=0)$. These two points correspond to spacetimes without a singularity.
@@ 3.5 General shape of the Master equation surfaces
In a generic case a solution to the Master equation defines a two-dimensional surface in the space of parameters $(a,b,c)$. The shape of this surface crucially depends on the parameter $\eta=A/C$ that is determined uniquely by the quantum conformal field theory in question. For a unitary CFT the parameter $\eta$ is positive and satisfies certain unitarity bounds that were studied in the literature. As we already explained this above in the present study we do not restrict ourselves to only unitary theories and thus prefer to explore all possible values of $\eta$ taking all possible (positive or negative) real numbers. In Figures 4-6 we present some typical situations with representative values of $\eta$.
The analysis shows that as $\eta$ varies in the interval $0<\eta<1$ the surface consists of two open shells. One of them lies in the quadrant of positive values of parameters $(a,b,c)$ and is bounded by plane $a+b+c=3$. So that for all points in this branch one has that $a+b+c>3$. The respective spacetime has regular integrated Riemann curvature, see (2.7).
The second open shell lies in the region of negative values of all parameters $(a,b,c)$. It contains a conical point that has coordinates $(0,0,0)$. It has three open infinite domains where the conditions (2.9) to have complete geodesics are satisfied. The respective spacetimes have infinite 4-volume near the singularity.
The entire surface is presented in Figure 4 for value $\eta=1/2$. Note that the points where all parameters vanish or where one of parameters is $1$ and the others vanish are not shown since they lie outside the continuous surface and represent isolated points in the parameters space.
For values $\eta>1$ the surface still has two open shells as in the case above. However, there also appear three closed identical shells each of them lying in a region where one of the parameters is negative and the two others are positive. A surface of this type is shown in Figure 5, where in particular the closed shell is shown in the region $(a<0,b>0,c>0)$. The closed shell then passes by the points $(a=0,b=1,c=0)$ and $(a=0,b=0,c=1)$ that now are not isolated points and are continuously related to other points in the closed shell component. Geometrically these three closed shells are identical due to the permutation symmetry in the parameter space $(a,b,c)$.
@@
In summary, we find that the configurations, that are similar to the classical Kasner solution, where one of the parameters is negative and the other two are positive, occur only for $\eta>1$ and $\eta<0$, and are absent for $0<\eta<1$. For $\eta>1$, the corresponding points lie on a compact surface, so the Kasner parameters remain bounded. In contrast, for $\eta<0$, the corresponding surface is non-compact, and the absolute values of the parameters are unbounded.
@@ 3.6 Solving the Master equation in terms of symmetric polynomials
The solution (3.12) has two branches depending on the sign. The expression under the square root is non-negative for two intervals: $e_{1}\geq 3$ and $e_{1}\leq\frac{1}{3}$. For $e_{1}=0$ in the positive sign branch $e^{-}_{2}=0$. This is the only point when $e_{2}$ vanishes. These values ($e_{1}=e_{2}=0$) imply that the parameters $(a,b,c)$ may take only the value $(0,0,0)$. This is a spacetime without any singularity.
@@
The bound (2.13) imposes an additional constraint on possible values of the symmetric polynomials. One finds that in the interval $e_{1}\geq 3$ the bound selects only the negative branch, $e^{-}_{2}$. In the interval $0\leq e_{1}\leq\frac{1}{3}$ both branches violate the bound while in the interval $e_{1}\leq 0$ the positive sign branch satisfies the bound. This is shown in Figure 7.
@@
Note, that this solution saturates the bound (2.13). It is possible only if all parameters $a,b,c$ are equal, $a=b=c=1$. We have already mentioned this solution above. It corresponds to a symmetric spacetime with a singularity at $t=0$.
@@
4 General form of anomalous stress tensor
In the preceding sections, we analyzed the trace part of the modified gravitational equations that incorporate the conformal anomaly. The full semiclassical gravitational equations (3.2), however, contain additional, tensorial, components that introduce further constraints. Geometrically, these constraints select a subspace of solutions within the surface ${\cal M}(a,b,c)=0$ considered earlier. To investigate their implications, one must first determine the general structure of the anomalous stress-energy tensor. In this section, we examine the complete set of equations for the stress-energy tensor $T_{\mu\nu}=\bigl(T_{00}(t),T_{11}(t),T_{22}(t),T_{33}(t)\bigr)$ of the quantum CFT and provide a full integration of these equations.